Ten less than twice Jose's age is 64. What is Jose's age?
Answer by gsmani_iyer(201)


Let Jose's age = x

So twice Jose's age = 2x
So 10 less than 2x means 2x-10
therefore, 2x-10 = 64 ......... Given
2x = 74
So x = 74/2 = 37.
Therefore Jose's age = 37 years.

I hope it is clear to you. Feel free to clear your doubts. I am a maths tutor.
All the best.